### v2

# Topological shape transform for thymus structures

Abstract
The Euler characteristic transform (ECT) is an emerging and powerful framework within topological data analysis for quantifying the geometry of shape. The applicability of ECT has been limited due to its sensitivity to noisy data. Here, we introduce SampEuler, a novel ECT-based shape descriptor designed to achieve enhanced robustness to perturbations. We provide a theoretical analysis establishing the stability of SampEuler and validate these properties empirically through pairwise similarity analyses on a benchmark dataset and showcase it on a thymus dataset. The thymus is a primary lymphoid organ that is essential for the maturation and selection of self-tolerant T cells, and within the thymus, thymic epithelial cells are organized in complex three-dimensional architectures, yet the principles governing their formation, functional organization, and remodeling during age-related involution remain poorly understood. Addressing these questions requires robust and informative shape descriptors capable of capturing subtle architectural changes across developmental stages. We develop and apply SampEuler to a newly generated two-dimensional imaging dataset of mouse thymi spanning multiple age groups, where SampEuler outperforms both persistent homology–based methods and deep learning models in detecting subtle, localized morphological differences associated with aging. To facilitate interpretation, we develop a vectorization and visualization framework for SampEuler, which preserves rich morphological information and enables identification of structural features that distinguish thymi across age groups. Collectively, our results demonstrate that SampEuler provides a robust and interpretable approach for quantifying thymic architecture and reveals age-dependent structural changes that offer new insights into thymic organization and involution.
Keywords: topological data analysis $|$ Euler characteristic transform $|$ thymus $|$ involution $|$ geometric morphometrics
Introduction
Morphological information has proven invaluable for understanding how structure relates to function in biological tissues, organs, and organisms [1, 2, 3]. Identifying and quantifying distinct shapes are central tasks in biology since geometric differences can indicate variations in developmental processes, disease states, or differences in biological function. The thymus is an essential organ of the adaptive immune system responsible for the maturation, selection, and differentiation of T cells [4]. Within the confines of the thymic microenvironment, T cells are instructed by thymic epithelial cells (TECs) and other stromal cells to establish a diverse, yet self-tolerant T cell receptor (TCR) repertoire - a process known as thymopoiesis [5]. With age, the thymus undergoes involution characterized by reduced total cell numbers, disruption of the epithelial network architecture, and diminished naive T cell output [6, 7]. Paralleling thymic involution and reduced thymopoietic activity, the TCR repertoire becomes constricted with age [8]. Together, this leads to an increased susceptibility to viral infections, heightened autoimmunity, and diminished surveillance of malignant cells due to the immune system’s diminishing capacity to mount effective responses to de-novo antigens [9, 10, 11]. Simultaneously, the thymus undergoes significant architectural remodeling, with both cortical TECs (cTECs) and medullary TECs (mTECs) exhibiting structural changes [2, 12]. To better understand these changes, precise mathematical descriptors are needed to quantify morphological changes at the cell level. A variety of shape descriptors have been developed for biological structures, ranging from deep learning frameworks that leverage large datasets to automatically learn shape representations [13], to techniques grounded in differential geometry that focus on curvature and other geometric invariants [14].
Recent advances in topological data analysis (TDA), a field dedicated to extracting geometric and topological insights from spatial data, have led to the introduction of persistent homology [15, 16] and to the development of the Euler Characteristic Transform (ECT) [17]. Extensions to ECT and refined transforms have been devised to capture aspects of shape analysis [18, 19, 20], and theoretical studies have established important properties of the ECT, including its invertibility [21, 22, 23] and the existence of stability bounds [24, 25]. The invertibility of ECT means that it captures all geometric and topological information of the input shape and is highly effective at capturing morphological differences between shapes. Despite the stability results, the ECT remains extremely sensitive to small perturbations and misalignments in the input shape, undermining its practical robustness. Although several alignment procedures have been proposed to address these issues [26, 27, 19], they often come at a high computational cost. Recent theoretical advances [22] have effectively resolved the sensitivity to misalignments, but the isometry invariant algorithms they have inspired [28] suffer from unsatisfactory levels of information loss.
Motivated by the limitations of existing isometry invariant algorithms, we introduce SampEuler, a discrete construction based on the ECT, which comes with theoretical guarantees about the information it captures and its stability. Building on SampEuler, we introduce a vectorization and visualization algorithm for the ECT. We first apply SampEuler and its vectorization to synthetic tree structures and show that they outperform current topological shape transforms. We then benchmark against conventional shape analysis methods and TDA methods using the MPEG7 dataset.
Finally, we apply SampEuler to a dataset of equally sized quadrants, extracted from binary images of mouse thymi across different age groups, to characterize age-related structural changes. In our classification study, we benchmark our methods against state-of-the-art TDA, including various vectorizations of persistent homology, as well as deep learning methods, demonstrating that SampEuler achieves higher classification accuracy in significantly less computation time. We interpret the classification results based on the vectorization of SampEuler in relation to underlying thymic structural differences, thereby providing valuable biological insights. Moreover, we compare the analysis of these quadrants based on shape descriptors with T-cell distribution analysis to study the relationship between the impact of ageing on T-cell distribution and thymic architecture.
Description of Datasets
We evaluate SampEuler on three datasets: synthetic structures for the method validation, established shape benchmarks for comparison with existing approaches, and mouse thymus images for biological applications.
Synthetic Networks
We produce simple geometric simplicial complex examples as networks consisting of three edges of fixed lengths emanating from a common origin that are uniformly discretized into points connected by edges in order. We perturb the points and the edges that connect them by adding Gaussian noise. This creates a controlled setting within which to evaluate the performance of ECT-based methods under perfect alignment versus rotational perturbations. Two representations of two groups of simplices are shown in Figure 1A. Details of the dataset can be found in the SI Appendix.
MPEG-7 Shape Library
We use the published image dataset from previous shape descriptor studies [29, 30, 31, 32, 33, 34, 17, 28, 35, 36, 37, 38]. The dataset consists of 70 groups of 2D binary images, each group containing silhouettes of objects from the same category. For example, one group contains 20 images of horse silhouettes. Although all the underlying objects are horses and share similar geometry, they are not identical; in many cases, a 2D isometry—such as a slight rotation—is applied to the object before the image is captured so that the same side of the horse is photographed from different angles.
We preprocess the dataset by first padding each image with black pixels to achieve a consistent size of $1126$ pixels $\times 1126$ pixels. Next, we denoise the images by retaining only the largest connected component of white pixels—treating any additional white pixels as false positives and converting them to black. The processed images are then used for further analysis. The preprocessing details are included in the SI Appendix.
Mouse Thymus Dataset
Mouse thymi (n=7) were processed as described in “Materials and Methods” (see below). The boundaries of the thymic epithelial cell (TEC) network were segmented using MATLAB’s Image Processing Toolbox: multilevel thresholding (multithresh) and skeletonization (bwskel) were applied to the cytokeratin 8 (K8) and cytokeratin 14 (K14) TEC markers to generate 2D binary images for all 7 thymi and both markers (values of 1 where the K8/K14 staining is present, 0 where it is absent). A 200 x 200 pixel lattice was superimposed on the 2D images to generate subimages, for subsequent TDA.
Results
We introduce SampEuler to capture shape information in a way that is robust to noise and overcomes the sensitivity of ECT (Figure 1). We show that the proposed approach achieves higher accuracy with less computational cost than existing methods, including persistent homology vectorizations, ECT-based approaches, and deep learning models (Figure 2). Employing SampEuler on the images of mouse thymi, we achieve age classification based on the shape of local thymic structures (Figure 3). Comparing with the distribution of T cells, we identify relationships between thymic structure and cellular composition at different ages (Figure 4).
From Data to Topological Transforms
Within TDA, the ECT is an integral transform that has been extended in numerous ways and successfully applied to biological datasets [18, 20, 27, 39, 19]. Such data comes in the form of n-dimensional images or discretized boundaries of segmented structures; these can be represented as simplicial complexes comprising points, edges, faces, and higher-dimensional simplices. One of the simplest computable topological invariants, the Euler characteristic, is given by the alternating sum of the number of n-dimensional simplices in the simplicial complex $K$: $\chi(K)=\sum_{i=1}^{n}(-1)^{i}\#\text{ of $i$ dimensional simplices in }K,$ where $n$ is the maximal dimension of all simplices in $K$. To capture geometric information of the complex, Turner et al. proposed sweeping a (hyper)plane through the shape progressively in time ($t$) in a given direction ($v$) and computing the Euler characteristic ($\chi$), which outputs an integer-valued curve for the particular direction (Equation 0.1). By sweeping through all directions, the Euler characteristic transform (ECT) is the correspondence between directions and curves (Equation 0.2). The ECT is invertible, meaning one can reconstruct the original shape and thereby derive any geometric information captured by other shape descriptors [21, 22, 23].
$\text{ECC}_{K}^{v}:\mathbb{R}\rightarrow\mathbb{Z};t\mapsto\chi(K\cap\{x\in\mathbb{R}^{n}:x\cdot v\leq t\}).$ (0.1)
$\text{ECT}(K):S^{n-1}\times\mathbb{R}\rightarrow\mathbb{Z};(v,t)\mapsto\text{ECC}_{K}^{v}(t).$ (0.2)
Since the ECT is sensitive to the complex, DETECT ([28]) was proposed to smooth, normalize, and average across all $\text{ECC}_{K}^{v}$:
$\text{DETECT}(K):[-a,a]\rightarrow\mathbb{R};$ (0.3)
$c\mapsto\int_{S^{n-1}}\left(\int_{-a}^{c}\text{ECC}_{K}^{v}(t)-\left(\frac{1}{2a}\int_{-a}^{a}\text{ECC}_{K}^{v}(s)\text{d}s\right)\text{d}t\right)\text{d}v$ (0.4)
We apply ECT and DETECT to the synthetic networks. To visualize the discriminative power of each method, we embed the resulting shape descriptors in two dimensions using multidimensional scaling (MDS) [40, 41]. The standard ECT distinguishes two classes of perfectly aligned shapes, whereas DETECT fails (Figure 1B-C, left), illustrating the information loss introduced during the averaging step. Indeed, both complexes yield identical DETECT transforms.
When the shapes are perturbed randomly to simulate the misalignment in applications, both methods fail to distinguish the two classes of complexes (Figure 1B–C, right).
Introduction to SampEuler
To overcome the sensitivity of ECT to perturbations on input shapes and the information loss of DETECT, we propose to record only the output curves and not the direction $v$ in ECC. This shape transform is isometry-invariant. To achieve this mathematically, we consider the ECT pushforward measure of $K$, which can be thought of as the probability law of ECCs of $K$ for a given direction chosen uniformly at random. The ECT pushforward measure is proven to be injective up to rotations and reflections [22].
By convention, the shape of a given complex is the equivalence class of the complex up to rotations, reflections, and translations. We centre all complexes by translating their barycentres to the origin to filter the translation information. On these centred complexes, the ECT pushforward measure serves as a complete shape descriptor: two complexes have the same ECT pushforward measure if and only if they have the same shape. The ECT pushforward measure is a probability law on the function space; and therefore is hard to compute in practice. Based on the ECT pushforward measure, we propose an isometry-invariant shape descriptor, SampEuler. SampEuler samples curves from the ECT pushforward measure and outputs an empirical measure formed by the sampled curves. We show that the SampEuler converges to the ECT pushforward measure as the discretization is refined. Equipped with the Wasserstein distance [42], it induces a continuous map from the input complex to its SampEuler representation. We also introduce a vectorization and visualization method for SampEuler. For small datasets, the vectorized representation yields a low-dimensional feature space that supports effective training of standard machine-learning models and, when paired with interpretation techniques, allows direct geometric interpretation. For larger or geometrically complex datasets, using the full SampEuler representation preserves complete geometric information with theoretical guarantees, avoiding information loss from dimensionality reduction. Details and proofs of SampEuler convergence and continuity are included in the SI Appendix.
We apply SampEuler and its vectorization to the synthetic networks, visualizing the results using MDS (Figure 1D-E). In both cases, SampEuler and its vectorization correctly distinguish the two classes (see SI Appendix for more details), demonstrating that they are both isometry invariant and informative.
Figure 1: MOTIVATION FOR NEW PUSHFORWARD MEASURE on ECT. We apply ECT, DETECT, SampEuler, and vectorization of SampEuler to aligned toy simplicial complexes(Left) and randomly rotated simplicial complexes (right).(A) Representations of simplicial complexes from the two classes. Each class shares the same base complex, in other words, the same shape, generated by attaching one end of three edges of fixed lengths with 100 vertices along each edge. We attach one end to the origin and randomly rotate the other end around the origin. For each sample, we add Gaussian noise at points along each edge. In the aligned case, we use the resulting complexes directly. In the rotated case, we apply individual random rotations centred at the origin to the whole complex around the origin. (B) Two-dimensional MDS embedding (a projection that preserves pairwise distances between shapes) of ECT shape descriptors. The results of two distinct classes can be separated when all complexes are aligned. However, in the rotated case, ECT results are heavily affected by the rotations and fail to distinguish two classes. (C) MDS embedding of DETECT shape descriptors. DETECT fails to distinguish two classes of shapes in both experiments due to information loss. (D) MDS embedding of SampEuler shape descriptors. In both experiments, the two classes are separated correctly. (E) MDS embedding of vectorized SampEuler shape descriptors. In both experiments, the two classes are separated correctly.
SampEuler Outperforms Topological Statistics
We apply SampEuler and its vectorization with Support Vector Machine (SVM) to classify the MPEG-7 dataset and benchmark against other methods. Details of image preprocessing and the experiments can be found in Materials and Methods.
We benchmark SampEuler against standard TDA methods that use the 10-class subset of MPEG-7: Persistence Scale Space Kernel [29], Persistence Weighted Gaussian Kernel [30], Sliced Wasserstein Kernel [31], Tangent Vector Representation with Gaussian Kernel [32], Persistence Fisher Kernel [33], and QUPID [34]. As shown in Figure 2B, SampEuler outperforms conventional TDA methods.
We run the classification task over the whole MPEG7 dataset to compare with other methods (Figure 2): Skeleton Paths [35], Bag of Contours [36], Persistence Image + Neural Network (PI+NN) [37], Persistence Codebook + Neural Network (PC+NN) [38]. SampEuler and its vectorization exhibit consistent performance; however, they are outperformed by dedicated 2D image recognition algorithms and TDA with deep learning approaches. We remark that topological methods offer more interpretability than deep learning, whose decision processes can be comparatively opaque.
Figure 2: CLASSIFICATION STUDY on MPEG7 DATASET. We compare SampEuler with conventional shape analysis methods using the MPEG7 dataset [43]. (A) Examples of the MPEG7 dataset images, one sample from each of the 10-class subset used in previous studies [33, 44]. (B) The barplot of the accuracy of each method in the classification task of the 10-class subset of MPEG7. We denote Persistence Scale Space Kernel [29] by $\mathcal{K}_{PSS}$, Persistence Weighted Gaussian Kernel [30] by $\mathcal{K}_{PWG}$, Sliced Wasserstein Kernel [31] by $\mathcal{K}_{SW}$, Tangent Vector Representation with Gaussian Kernel [32] by $\mathcal{K}_{TVRG}$, Persistence Fisher Kernel [33] by $\mathcal{K}_{PF}$. (C) The barplot of the accuracy of each method in the classification task of the whole MPEG7 dataset (70 classes). We use methods: Skeleton Paths [35], Bag of Contours [36], Persistence Image + Neural Network (PI+NN) [37], Persistence Codebook + Neural Network (PC+NN) [38], ECT, DETECT, SampEuler and vectorization of SampEuler.
To understand the effect of discretizing ECT and ECT pushforward in practice, we use the MPEG-7 dataset and vary the number of filtration values sampled along each curve and the total number of curves sampled, and perform classification tasks with SVM. Experimental details and full results can be found in SI Appendix. We report that if we sample more than 100 directions and 1000 points along each direction, we obtain high accuracy, and the additional information obtained for the MPEG7 dataset is limited. Moreover, leveraging the optimised implementation in Eucalc [45] enables efficient computation of these descriptors, making the above parameter choices practical in routine use.
SampEuler Quantifies Ageing Impact on Thymic Structure
To assess how ageing affects local architecture of TEC, we sub-sampled quadrants of size $200\times 200$ pixels in representative areas of the cortex (i.e. K8 segmentation mask) and medulla (K14 segmentation mask). We sampled 20 quadrants per thymic lobe, resulting in a total of 40 quadrants for young and 100 for old. They were used as input to a supervised, binary classification of young versus old (Figure 3A). For this resolution, we compute $360$ directions and $300$ points along each direction for all ECT-based methods.
We benchmark our proposed methods against the Persistence image [37] with the signed distance function, AlexNet [46], Visual Transformer [47], MobileNet V2 [48], EfficientNetB0 [49], ECT and DETECT to verify that they accurately quantify local thymic structure. Details of the experiments can be found in SI Appendix. The results are shown in Figure 3B-C. SampEuler and its vectorization give better accuracy with shorter run times, demonstrating their ability to efficiently detect age-related structural differences within the cortex or medulla.
To interpret geometric differences between age groups, we use the vectorization of SampEuler to visualize Euler characteristic curves across all directions. We apply the SHAP algorithm to identify which regions of this image are most important for classifying each age group. We repeat training 50 times, retaining trials with accuracy above 80% as valid. For each age group, we average SHAP values across test samples from all valid trials to obtain a final feature importance map (Figure 3D). These maps reveal the distribution of Euler characteristic curves that the classifier recognizes as typical for each age group: regions of high importance indicate where curves consistently lie, effectively summarizing each group’s typical shape signature.
In the cortical region (K8) of the thymus, both age groups exhibit an initial increase in Euler characteristic with increasing filtration value $t$, with a slightly steeper rise in young thymi. At higher values of $t$, older thymi show a markedly steeper decline in Euler characteristic and reach a lower minimum, whereas younger thymi decrease more gradually. Since the Euler characteristic for the images is computed from their corresponding 2D cubical complexes, we have that
$\text{Euler characteristic }=\#(\text{connected components})-\#(\text{loops}).$
Here, the number of loops means the dimension of the first homology group of the cubical complex or, intuitively, the number of holes in the image. Using this relationship, we interpret the observed differences as follows: the steeper initial increase in young thymi suggests a higher density of connected components; the lower final Euler characteristic in older thymi indicates more loops in the cortical structure; and the steeper decline in older thymi implies these loops are more densely distributed.
In the medulla (K14) of the thymus, the major trends of both age groups are similar. However, we observe greater variation in the geometry of old thymi compared to young thymi. Old quadrants show more diverse patterns in their Euler characteristic curves, with a higher proportion exhibiting increased numbers of connected components. In particular, we use other methods to confirm these interpretations (see SI Appendix).
We also analyzed how morphological and cellular composition differences between age groups vary with cortical depth (Figure 3E). For each thymus, we defined a depth kernel ranging from 0 at the outer capsule to 1 at the cortico-medullary junction (CMJ). At each depth, we used energy distance [50] to quantify differences between age groups based on pairwise SampEuler distances (capturing structural differences) and T-cell enrichment ratios (capturing cellular composition across whole tiles; see SI Appendix for definitions). The results show that both morphological and cellular differences between age groups occur near the capsule and diminish towards the CMJ, highlighting coincident changes between cTEC morphology and cellular composition with age.
Figure 3: CLASSIFICATION STUDY ON SELECTED THYMIC QUADRANTS. We segment masked images of thymic epithelium into quadrants of size $200\times 200$ pixels. Then, we classify them into old and young groups separately for K8 and K14 markers. (A) Example images of thymic quadrants of each age group and each marker. (B) Barplots of the classification accuracies of each method. The top plot contains accuracies for classifying K8 quadrants, and the bottom plot contains accuracies for classifying K14 quadrants. The error bars show the standard deviations for repeating the classification 50 times. (C) Barplots of the classification runtimes of each method. The top plot contains runtimes in minutes for classifying K8 quadrants, and the bottom plot contains runtimes in minutes for classifying K14 quadrants. (D) We use the SampEuler vectorization as input features for classification training. We repeat the train-test process 50 times and keep trials with over $80\%$ accuracy. For trials with accuracy over $80\%$, we compute the SHAP value for each feature. For each marker and age group, we average across all test samples of the age group and then across all valid trials. The plot shows average SHAP values for each feature as feature importances towards classifying the sample as a given age group. (E) Depth-dependent comparisons between thymic quadrants from different age groups. For both plots, the x-axis shows the normalized depth of the thymus, where 0 corresponds to quadrants at the capsule and 1 to quadrants at the cortex–medulla junction (CMJ). In the left plot, the y-axis shows the energy distance [50] between age groups based on T-cell enrichment ratios. In the right plot, the y-axis shows the energy distance between age groups based on pairwise SampEuler distances.
Comparing Cell Distribution with Shape
In addition to TEC segmentation, we annotated the hematopoietic and stromal cell types for this dataset. To compare the impact of ageing on both cell distributions and TEC architecture, we perform k-medoids clustering [51] separately based on cell enrichment ratios (see SI Appendix for definitions) and SampEulers.
Age-related thymic involution impacts medullary and cortical TEC types differentially. While both cell populations decrease in absolute terms, the relative representation of cTEC to mTEC increases with age [7]. cTECs are known to mediate key biological processes like lineage commitment and positive T cell selection [52]. These processes are essential for hematopoietic T cell precursors to a) adopt a T cell lineage, and later in development to b) probe their ability to bind self-peptide on MHC molecules. Therefore, we decided to focus on the cortical region for the subsequent analyses and investigate morphological changes in the cortical region along with compositional changes in proximal, abundant hematopoietic cell types. For the composition definition we limited this analysis to the most abundant and biologically informative hematopoietic cell types in the cortical compartment: double-negative type 3 (DN3) T cells, pre (positive) selection double-positive T cells (Presel DP), post (positive) selection double positive T cells (CD69-) (Postsel DP, CD69-), and post (positive) selection double positive T cells (CD69+) (Postsel DP, CD69+). Full phenotype marker definitions are provided in SI Appendix. Furthermore, we restricted the compositional analysis to those thymocytes that lie within $5\mu\text{m}$ distance to the K8 segmentation mask. We clustered all cortical quadrants from both age groups together so that cluster labels are comparable across age.
To allow faster computation of the pairwise Wasserstein distance matrix of all cortex quadrants, we use the sliced Wasserstein distance algorithm [53] with $50$ slices to approximate the true Wasserstein distance.
[…]
Discussion
In this paper, we introduced and implemented SampEuler and its vectorization as two novel topological shape descriptors. We applied them to a benchmarking shape image dataset and a mouse thymus dataset. The results show that both novel shape descriptors are robust to rigid transformations, resolving the high sensitivity issue of ECT [17]. Moreover, SampEuler can extract shape information effectively to resolve subtle shape differences with high accuracy.
We demonstrated the effectiveness of SampEuler and its vectorization on synthetic three-edge tree simplices, highlighting their capacity to capture and emphasize purely shape-based information. Benchmarking on the MPEG7 dataset confirmed that SampEuler achieves higher classification accuracy than conventional TDA methods based on persistent homology. While dedicated 2D image recognition algorithms and TDA with deep learning approaches achieve higher accuracy, SampEuler provides a favorable balance between performance and interpretability. Their application to the mouse thymus dataset underscored the strength of SampEuler in the study of complex biological structures. The random forest classifiers trained on SampEuler and the vectorization of SampEuler achieved competitive accuracies ($70\%$) for age classification relying on a limited set of training data (i.e. cell shapes) at notably lower computational cost than deep learning models. This shows that a) a quantifiable change in cell morphology occurs with age for both cTEC and mTEC and b) this shape change can inform the age-category of the underlying sample. While this holds true for the two ages considered, a more granular approach (extending over more time points) is required to confirm its validity across a broader life trajectory. Applying the SHAP algorithm to random forest classifiers trained on SampEuler vectorizations enabled a mathematical characterization of cell shape. Although classification reveals detectable shape differences between age groups, these age-related changes appear to affect different morphological subtypes in parallel: each shape cluster undergoes its own characteristic shift with age, but the relative distinctions between clusters are preserved. Consequently, the overall spatial organization of shape clusters remains largely intact (Figure 4A i and iii). However, when considering cell composition clustering, a spatial shift becomes apparent with age. In young thymi, Cluster A is confined to the sub-capsular region, whereas in old thymi, this same compositional pattern extends deep into the cortical compartment (Figure 4A ii vs. iv). This spatial reorganization could arise from several age-related factors, such as loss of intra-thymic cytokine gradients that normally maintain youthful regional organization, or failure in the provision of lympho-stromal cross-talk necessary for double-negative T cell progression. However, additional experimental validation is required to confirm these hypotheses.
Building on these positive findings, we propose that future work investigate the theoretical properties of the ECT pushforward measure in general simplicial complex settings, as well as develop formal results on the information captured by the vectorization of SampEuler. While the present study focuses on 2D image masks, the pipeline extends naturally to simplicial complexes in both 2D and 3D. In particular, applying the framework to 3D images would better reflect the underlying biological structures and further enhance the quality of the analysis pipeline introduced in this work.
Materials and Methods
MPEG-7 Dataset
The MPEG-7 shape dataset was used for benchmark shape analysis methods. We applied algorithms to fill in the holes within the objects and retained only the largest connected component inside the image. We then normalized the image sizes to be equal across all classes (See SI Appendix for details).
Thymus Experimental Procedure
[…]
Thymus Image Processing and Cell Annotation
[…]
Euler Characteristic Transform (ECT)
ECT computations for grayscale images were performed using the Eucalc package (https://github.com/HugoPasse/Eucalc). The algorithm normalizes images to a size of 1 unit by 1 unit. We chose the range for the parameter of sublevel sets as (1.5,1.5)(-1.5,1.5) to allow full inclusion of the image along all possible directions. The number of directions and sublevel set parameters were tuned for each dataset; an ablation study examining parameter sensitivity is provided in SI Appendix.
Ethics Statement
Both the caring for the mice and the experiments were performed in compliance with permissions of the Cantonal Veterinary Office of Basel-Stadt, Switzerland, and in accordance with Swiss federal regulations.
Data Availability
Code for generating synthetic networks, preprocessing MPEG-7 images, and computing ECT-based methods for simplicial complexes is available at https://github.com/reddevil0623/Shape-transform-descriptor-for-thymus-structures (DOI: https://doi.org/10.5281/zenodo.18614535). Thymus segmentation mask quadrants, total masks, and cell composition data for each quadrant are also available at this repository. The original MPEG-7 dataset is publicly available from https://dabi.temple.edu/external/shape/MPEG7/dataset.html. Raw cell-level thymus data are available from the corresponding author upon reasonable request.
Acknowledgments
HAH, HMB, VL are grateful for the support provided by the UK Centre for Topological Data Analysis Engineering and Physical Sciences Research Council (EPSRC) grant EP/R018472/1 and EPSRC EP/Z531224/1. HAH gratefully acknowledges funding from the Royal Society RGF/EA/201074, UF150238 and EPSRC EP/Y028872/1 and EP/Z531224/1. HAH and HY acknowledge funding by the Leverhulme Trust Prize PLP-2020-252. HY is grateful for the support of the Leathersellers’ Foundation. HMB acknowledges support provided by the Mark Foundation for Cancer Research. AT and GAH are supported by the Wellcome Trust (Wellcome Collaborative Award SynThy, 211944/Z/18/Z), the Swiss National Science Foundation (310030B_138655 and 310030_215113), and the DiGeorge Syndrome Research Fund at Stanford University, generously established by an anonymous donor. LT gratefully acknowledges funding from Foundation Immune Engineering for Global Child and Adolescent Health.
Author Contributions
H.A.H., G.A.H. and H.M.B. designed research; H.Y., A.T., L.T. and V.L. performed research; H.Y., V.L., H.A.H., and H.M.B. developed new analytic tools; S.Z. performed the experiments; H.Y., A.T. and L.T. analyzed data; and H.Y., V.L., A.T., L.T., G.A.H., H.A.H., and H.M.B. wrote the paper.
Competing Interests
The authors declare no competing interests.
Supporting Information Appendix
Appendix S1 Data
We use three sources of data: synthetic network data, published MPEG-7 Core Experiment Shape-Matching dataset (MPEG-7) [57], and experimental image data of mouse thymi of different age groups. We explain the details of each dataset in turn: first the synthetic network dataset, then the preprocessing performed on the MPEG-7 dataset. Finally, we describe the thymus experiment and image processing details.
S1A Synthetic Networks
We generate a class of synthetic networks as follows. We construct each sample by first generating three edges of lengths $2$, 33, and $4$ with their common endpoint fixed at the origin $(0,0)$. Each edge is discretized into $n=100$ equally‐spaced vertices (including the origin). For each edge, we select a rotation angle $\varphi\sim\mathrm{Uniform}(0,2\pi)$ and rotate all vertices about $(0,0)$ by $\varphi$. For each non-origin vertex, we independently sample Gaussian noise
$(\varepsilon_{x},\varepsilon_{y})\sim\mathcal{N}\bigl(0,\,\sigma^{2}\bigr),\quad\sigma=0.02,$
and obtain the perturbed vertex position by adding this noise to the rotated coordinates. Connectivity (line segments between consecutive edge points and from the origin to the first point of each edge) is preserved. We repeat this process $k=20$ times to generate $20$ structures for each shape class. We generate two different shape classes for comparison by choosing different angles of the three edges at the start.
S1B MPEG-7 Core Experiment Shape-Matching dataset
We use the standard MPEG-7 Core Experiment Shape-Matching (MPEG-7) dataset [57], which contains 1400 binary images of silhouettes. The images are split into 70 categories, each containing 20 images based on the same types of objects (e.g., apples, horses, etc.).
We preprocess the images by first expanding the white connected component by 1 pixel in all directions to close small gaps and filter to keep only the largest connected component. We then fill in the holes of the filtered images using [58]. To normalize across the dataset, each remaining white component is then rescaled to a fixed target area, and the image is padded with black pixels until it matches a common output size. Finally, we translate the white component so that the centroid of its white pixels coincides with the image centre. The preprocessing code is available at: https://github.com/reddevil0623/Shape-transform-descriptor-for-thymus-structures
Figure S1: A: Pipeline for generating a class of tree samples. We first generate three edges of lengths 2, 3, and 4 from the origin, randomly oriented. We discretize each edge by sampling 100 equally spaced points along it and connecting consecutive points with edges. For each sample in the class, we randomly rotate the tree around the origin and then add coordinate-wise Gaussian noise of $\sigma=0.02$ to each vertex. We keep the original connecting order of the edges. B: Preprocessing pipeline of MPEG-7 Core Experiment Shape-Matching dataset. For each image, we first expand all white connected components by 1 pixel in all directions to close small gaps and filter to keep the largest connected component. We then fill in all the holes for the filtered image using [58]. We pad black pixels around the original image to normalize all images to the same size, with the centroid of the connected component at the center of each padded image. C: Four example image quadrants of the mouse thymus dataset, each of size 200 $\times$ 200 pixels. The pixels are of side length $0.5\text{ }\mu\text{ m}$. From left to right, top to bottom, they are from young-cortex, old-cortex, young-medulla, and old-medulla.
S1C Thymus Experimental Procedure, Image Processing, and Cell Annotation
[…]
Appendix S2 Analysis Methods
In this section, we introduce the necessary background on topological transforms and theoretical extensions in the first section, and then in the second section, we describe other analysis methods used in this study.
S2A Topological Transforms
We recall the Euler characteristic transform (ECT) and then present a theoretical extension (Theorem S6 and Theorem S7). Next we introduce SampEuler and present desirable theoretical and practical properties of this descriptor for analysing shape datasets (Theorem S9).
The Euler Characteristic Transform
The Euler characteristic transform (ECT) is one of the foundational tools within the Topological Data Analysis community [27, 28, 19, 20].
In this section, we introduce the basic concepts of ECT for geometric simplicial complexes in order to describe our analysis methods. See [59] and [22] for a more general setting of ECT.
Definition S1.
A geometric $k$-simplex $K_{1}$ is the convex hull of $k+1$ affinely independent points $v_{0},v_{1},\dots,v_{k}$ and denoted by $v_{0},\dots,v_{k}]$. For any $\{u_{0},\dots,u_{j}\}\subset\{v_{0},\dots,v_{k}\}$, we call $u_{0},\dots,u_{j}]$ a face of $v_{0},v_{1},\dots,v_{k}$. The dimension of $K_{1}$ is $\dim(K_{1})=k$.
A geometric simplicial complex $K$ is a finite collection of geometric simplices $\{K_{1},\dots,K_{n}\}$ such that
  1. (i)
    For every $K_{i}\in K$, every face of $K_{i}$ is also an element of $K$;
  2. (ii)
    For any $K_{i},K_{j}\in K$, either $K_{i}\cap K_{j}=\emptyset$ or $K_{i}\cap K_{j}$ is a face of both $K_{i}$ and $K_{j}$.
The dimension of $K$ is defined as $\max\{\dim(K_{i}):i\in\{1\dots n\}\}$.
Definition S2.
For a geometric $k$-simplex $K_{1}$, the Euler characteristic of $K_{1}$, denoted by $\chi(K_{1})$, is defined to be $\chi(K_{1}):=(-1)^{k}$. For a geometric simplicial complex $K$ consisting of the finite collection $\{K_{1},\dots,K_{n}\}$, the Euler characteristic of $K$ is defined as $\chi(K)=\sum_{i=1}^{n}\chi(K_{i})$.
Example 1.
For any discretization $K$ of the unit circle, there are $n$ vertices (0-simplices) and $n$ edges ($1$-simplices) forming a closed cycle for some integer $n$. Therefore, the Euler characteristic is $\chi(K)=n-n=0$ for any such discretization $K$.
Definition S3.
For geometric simplicial complex $K$ consists of simplices $K_{1},\dots,K_{n}$ of dimension $d$, the Euler characteristic transform of $K$, denoted by $\text{ECT}(K):S^{d-1}\times{\mathbb{R}}\rightarrow{\mathbb{Z}}$ is defined as:
$\text{ECT}(K)(v,t)=\chi(K^{\langle v,\cdot\rangle\leq t}),$
where $K^{\langle v,\cdot\rangle\leq t}=\{K_{i}:\langle v,x\rangle\leq t,\forall x\in K_{i}\}$. Fixing direction $v$, the curve $\text{ECT}(K)(v,-):{\mathbb{R}}\rightarrow{\mathbb{Z}}$ is the Euler characteristic curve (ECC) of $K$ along direction $v$.
We now briefly recall the notion of persistent homology and persistence diagrams; see [17, 22] for details in the context of topological transforms.
Given a topological space $X$, its $k$-th homology (with coefficients in a field) is a vector space whose dimension counts the number of independent $k$-dimensional “holes” of $X$. By 0-dimensional holes we mean connected components, by $1$-dimensional holes we mean cycles or loops, and by $2$-dimensional holes we mean enclosed voids. These $k$-dimensional holes are often called homological features of $X$.
The idea of persistent homology is to track how these homological features appear and disappear as the simplicial complex grows through a one-parameter filtration. Given a sequence of complexes $(K_{t})_{t\in{\mathbb{R}}}$ such that $K_{t}\subset K_{r}$ for all $t\leq r$, each homological feature is born at some parameter value $a\in{\mathbb{R}}$ and dies at some $b>a$. The pair $(a,b)$ records the birth and death of that feature. The multiset of all such pairs $\{(a_{i},b_{i})\}$ is called the persistence diagram. Points far from the diagonal $\{a=b\}$ correspond to long-lived, and hence typically more significant, topological features, while points close to the diagonal represent short-lived features that are often attributed to noise.
In the context of topological transforms, for a fixed direction $v\in S^{d-1}$, the height filtration $K^{\langle v,\cdot\rangle\leq t}$ gives rise to a one-parameter family of subcomplexes of $K$ from which we can compute a persistence diagram for each homological degree.
Definition S4.
For geometric simplicial complex $K$, the persistent homology transform of $K$ denoted by $\text{PHT}(K):S^{d-1}\rightarrow\mathrm{Dgm}^{d}$ is defined as:
$\text{PHT}(K)(v)=(\text{PH}_{0}(K,v),\dots,\text{PH}_{d-1}(K,v))$
where $\mathrm{Dgm}$ is the space of persistence diagrams, and $\text{PH}_{i}(K,v)$ is the $i^{\text{th}}$ persistent homology of $K$ induced by the height filtration along direction $v$.
We use $\operatorname{CF}({\mathbb{R}})$ to denote the space of all piecewise constant functions from $\mathbb{R}$} to $\mathbb{Z}$} as an integer combination of half-closed and half-open intervals. The ECCs resulting from ECTs are within $\operatorname{CF}({\mathbb{R}})$.
Theorem S1 (Continuity and Hausdorffness [22]).
Given a geometric simplicial $K\subset{\mathbb{R}}^{d}$, if the ECT of $K$ is viewed as $\text{ECT}(K):S^{d-1}\rightarrow\operatorname{CF}({\mathbb{R}})$, a map from the space of directions to the space of ECCs, then $\text{ECT}(K)$ is Lipschitz continuous with normal metric on $S^{d-1}$ and $L_{p}$ metric on $\operatorname{CF}({\mathbb{R}})$ with $1\leq p<\infty$. Furthermore, the space $\operatorname{CF}({\mathbb{R}})$ is a Hausdorff space.
We fix $p=1$ and adopt the $L_{1}$ metric for $\operatorname{CF}({\mathbb{R}})$ below, but the results hold true for any $1\leq p<\infty$ unless stated explicitly.
Theorem S2 (Injectivity theorem [22, 23]).
Consider two geometric simplicial complexes $K$ and $L$. We have $\text{ECT}(K)=\text{ECT}(L)$ if and only if $K=L$.
The injectivity result shows that ECT can capture all geometric and topological information of the input complex. However, some differences between various complexes may not be of interest in analysis. For instance, a nonsymmetric complex $K$ before and after applying a nontrivial rotation would generate two different ECT results.
As suggested by [60], the shape of a geometric simplicial complex $K$ is the equivalence class of $K$ up to rigid motions, i.e., translations, rotations, and reflections. Centring the complex $K$ such that the sum of all vertices of $K$ is at the origin enables rapid alignment of shapes up to translations. Throughout this paper, we assume the input complex is centred. However, no computationally cheap alignments can be performed for isometries.
Definition S5.
Given two geometric simplicial complexes $K$ and $L$, we write $K\sim L$ if and only if there exists a rigid motion $\varphi$ such that $\varphi(K)=L$. The relation $\sim$ is an equivalence relation, and we define the shape class of $K$ as the equivalence class $[K]$.
We use the following definitions from [22].
Definition S6.
A geometric simplicial complex $K$ in ${\mathbb{R}}^{d}$ with vertex set $X$ is strongly generic if:
  1. (i)
    ECCs are distinct for all directions $v\in S^{d-1}$,
  2. (ii)
    the vertex set $X$ is in general position, i.e., no subset of $k$ points in $X$ lie on a (k2)(k-2)-dimension affine subspace, for $k=2,3,\dots,d+1$.
Definition S7.
For strongly generic geometric simplicial complexes $K\subset{\mathbb{R}}^{d}$, the ECT pushforward measure of $K$ is $\text{ECT}(K)_{*}\mu:\mathcal{B}(\operatorname{CF}({\mathbb{R}}))\rightarrow{\mathbb{R}}$, such that $\text{ECT}(K)_{*}\mu(E)=\mu(\{v\in S^{d-1}:\text{ECT}(K)(v,-)\in E\})$, where $\mu$ is the uniform measure on the sphere and $\mathcal{B}(\operatorname{CF}({\mathbb{R}}))$ is the Borel $\sigma$-algebra on the space of ECCs.
Theorem S3 (Theorem 6.7 of [22]).
Let $K$ and $L$ be strongly generic geometric simplicial complexes in ${\mathbb{R}}^{d}$. Let $\mu$ be the Lebesgue measure on $S^{d-1}$. If the pushforward measures $\text{ECT}(K)_{*}(\mu)=\text{ECT}(L)_{*}(\mu)$, then there is some $\varphi\in O(d)$ such that $L=\varphi(K)$, i.e. $L$ and $K$ are identical up to a combination of rotations and reflections.
This shows that the ECT pushforward measure characterizes the shape classes of the complexes completely. Based on the ECT pushforward measure, Marsh et al. defined the Detecting Temporal shape changes with the Euler Characteristic Transform (DETECT) algorithm and applied it successfully to biological datasets [28].
Definition S8.
Given a geometric simplicial complex $K\subset{\mathbb{R}}^{d}$, and suppose that the norms of all vertices are bounded within interval [a,a][-a,a]. The DETECT curve of $K$ is defined as:
$\text{DETECT}(K):{\mathbb{R}}\rightarrow{\mathbb{R}}$
$t\rightarrow\int_{S^{d-1}}\int_{-a}^{t}\text{ECT}(K)(v,s)-\overline{\text{ECT}(K)(v,-)}\,\mathrm{d}s\,\mathrm{d}v$
where $\overline{\text{ECT}(K)(v,-)}=\int_{-a}^{a}\text{ECT}(K)(v,s)\,\mathrm{d}s$.
Stability properties are important for data analysis. We first introduce some metrics and prove the stability property for the ECT pushforward measure in Theorem S7.
Definition S9.
Let $(M,d)$ be a metric space and $\mathcal{P}_{p}(M)$ be the collection of all probability measures $P$ such that $\int_{M}d(x,x_{0})^{p}\text{d}P(x)<\infty$ for some $x_{0}$. The $p^{th}$ Wasserstein distance for $P,Q\in\mathcal{P}_{p}(M)$ is defined as:
$W_{p}(P,Q):=\left(\inf_{\gamma\in\Gamma(P,Q)}\left\{\int_{M\times M}d(x,y)^{p}\text{d}\gamma(x,y)\right\}\right)^{1/p}$
where $\Gamma(P,Q)$ is the set of all measures $\mu$ on $M\times M$ such that $\mu(A\times M)=P(A)$ and $\mu(M\times B)=Q(B)$.
Let $\Delta=\{(x,x):x\in\mathbb{R}\}$ denote the diagonal in $\mathbb{R}^{2}$. For a point $(b,d)\in\mathbb{R}^{2}$, define its projection onto the diagonal as
$\pi_{\Delta}(b,d)=\left(\frac{b+d}{2},\frac{b+d}{2}\right).$
Definition S10 (Matching).
Given two persistence diagrams $D$ and $D^{\prime}$, a matching between $D$ and $D^{\prime}$ is a map
$\gamma:D\cup\Delta\to D^{\prime}\cup\Delta$
where the diagonal $\Delta$ is taken with infinite multiplicity, such that:
  1. (i)
    Each point $x\in D$ is either matched to a point $\gamma(x)\in D^{\prime}$, or matched to its projection $\pi_{\Delta}(x)\in\Delta$;
  2. (ii)
    Each point $y\in D^{\prime}$ is either matched to some point in $D$, or matched to its projection $\pi_{\Delta}(y)\in\Delta$.
We denote the set of all such matchings by $\Gamma(D,D^{\prime})$.
Definition S11.
[61] For $p\in[1,\infty]$, the $p$-Wasserstein distance between persistence diagrams $D$ and $D^{\prime}$ is defined as
$W_{p}(D,D^{\prime})=\left(\inf_{\gamma\in\Gamma(D,D^{\prime})}\sum_{x\in D\cup\Delta}\|x-\gamma(x)\|_{p}^{p}\right)^{1/p},$
where $\|\cdot\|_{p}$ denotes the $\ell^{p}$-norm on $\mathbb{R}^{2}$.
If $p=\infty$, the Wasserstein distance is defined as:
$W_{\infty}(D,D^{\prime})=\inf_{\gamma\in\Gamma(D,D^{\prime})}\max_{x\in D\cup\Delta}\|x-\gamma(x)\|_{\infty}.$
Recall the following definition taken from [22].
Definition S12.
Given two geometric simplicial complexes $K$ and $L$, the ECT metric between $K$ and $L$ is defined as:
$d_{\text{ECT}}^{1}(K,L)=\int_{S^{d-1}}\|ECT(K,v)-ECT(L,v)\|_{1}\,\mathrm{d}v,\text{ for }1\leq p<\infty.$
The $\infty$-PHT distance in degree $i>0$ between $K$ and $L$ is defined as:
$d_{\infty}^{\text{PHT}^{i}}(K,L)=\max_{v\in S^{d-1}}W_{\infty}(\text{Dgm}_{i}(K,v),\text{Dgm}_{i}(L,v)).$
where $\text{Dgm}_{i}(K,v)$ and $\text{Dgm}_{i}(L,v)$ are the $i^{\text{th}}$ persistence diagrams of $K$ and $L$ induced by height filtration along direction $v$.
Theorem S4 (Proposition 2 of [62]).
For geometric simplicial complexes $K$ and $L$ in ${\mathbb{R}}^{d}$, we have:
$\|\text{ECT}(K,v)-\text{ECT}(L,v)\|_{1}\leq\sum_{k=0}^{d-1}2n_{k}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v)),$
where $\text{Dgm}_{k}(K,v)$ and $\text{Dgm}_{k}(L,v)$ are the $k^{\text{th}}$ persistence diagrams of $K$ and $L$ induced by height filtration along direction $v$, and $n_{k}$ is the maximum number of points in $\text{Dgm}_{k}(K,v)$ or $\text{Dgm}_{k}(L,v)$.
Definition S13.
Let $K,L\subset{\mathbb{R}}^{d}$ be geometric simplicial complexes and let $\varphi:K\rightarrow L$ and $\psi:L\rightarrow K$ be a homotopy equivalence of $K$ and $L$. Let $H_{K}:K\times I\rightarrow K$ and $H_{L}:L\times I\rightarrow L$ be homotopies from $\psi\circ\varphi$ to $\text{Id}_{K}$ and from $\varphi\circ\psi$ to $\text{Id}_{L}$ respectively. We denote by $\|\cdot\|_{{\mathbb{R}}^{d}}^{2}$ the normal Euclidean 2-norm. If there exists $\epsilon$ strictly positive such that $\|x-\varphi(x)\|^{2}_{{\mathbb{R}}^{d}}\leq\epsilon$ and $\|y-\psi(y)\|^{2}_{{\mathbb{R}}^{d}}\leq\epsilon$ for all $x\in K$ and $y\in L$, as well as $\|x-H_{K}(x,s)\|^{2}_{{\mathbb{R}}_{d}}\leq 2\epsilon$ and $\|y-H_{L}(y,s)\|^{2}_{{\mathbb{R}}_{d}}\leq 2\epsilon$ for all $x\in K,y\in L,s\in I$, then we say $K$ and $L$ are $\epsilon$-controlled homotopic.
Theorem S5 (Lemma 4.5 and Theorem 4.11 of [63]).
Let $K,L\subset{\mathbb{R}}^{d}$ be geometric simplicial complexes that are $\epsilon$-controlled homotopic. Then we have:
$\max_{i}d_{\infty}^{\text{PHT}^{i}}(K,L)\leq\epsilon.$
Theorem S6.
Let $K,L\subset{\mathbb{R}}^{d}$ be geometric simplicial complexes that are $\epsilon$-controlled homotopic, and let $n_{k}$ be the maximum number of bars in $\text{Dgm}_{k}(K,v)$ or $\text{Dgm}_{k}(L,v)$ for all $v\in S^{d-1}$, and let $n=\sum_{k=0}^{d-1}n_{k}$. Then we have:
$d_{\text{ECT}}^{1}(K,L)\leq 2dn\epsilon.$
Proof.
By the definition of the ECT metric we have:
$d_{\text{ECT}}^{1}(K,L)$ $=\int_{S^{d-1}}\|\text{ECT}(K,v)-\text{ECT}(L,v)\|_{1}\,\mathrm{d}v$
$\leq\int_{S^{d-1}}\sum_{k=0}^{d-1}2n_{k}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v))\,\mathrm{d}v$
$=\sum_{k=0}^{d-1}2n_{k}\int_{S^{d-1}}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v))\,\mathrm{d}v$
$\leq\sum_{k=0}^{d-1}2n_{k}\max_{v\in S^{d-1}}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v))$
$=2\sum_{k=0}^{d-1}n_{k}d_{\infty}^{\text{PHT}^{k}}(K,L)\leq 2dn\epsilon.$
Theorem S7.
The ECT pushforward measure equipped with the Wasserstein distance $W_{1}$ is Lipschitz continuous for homotopic equivalent complexes.
Proof.
We know that ECT is Lipschitz continuous when equipped with $d_{\text{ECT}}^{1}$ metric from Theorem S6. It is sufficient to show that for two complexes $K,L\subset{\mathbb{R}}^{d}$, the Wasserstein distance between two pushforward measures is bounded by the ECT metric. Let $\|f-g\|_{1}=\int|f(t)-g(t)|\,\mathrm{d}t$ be the $L_{1}$ function metric. We have:
$W_{1}(\text{ECT}(K)_{*}\mu$ $,\text{ECT}(L)_{*}\mu)=\inf_{\gamma\in\Gamma(\text{ECT}(K)_{*}\mu,\text{ECT}(L)_{*}\mu)}\left(\int_{\text{CF}({\mathbb{R}})\times\text{CF}({\mathbb{R}})}\|f-g\|_{1}\,\mathrm{d}\gamma(f,g)\right)$
$=\inf_{\gamma\in\Gamma(\text{ECT}(K)_{*}\mu,\text{ECT}(L)_{*}\mu)}\left(\int_{\text{Im}(\text{ECT}(K))\times\text{Im}(\text{ECT}(L))}\|f-g\|_{1}\,\mathrm{d}\gamma(f,g)\right).$
Let $\nu_{r}(K,L)$ be the pushforward of the uniform measure on the sphere by the map $\text{ECT}(K,-)\times\text{ECT}(L,r(-)):S^{d-1}\rightarrow\operatorname{CF}({\mathbb{R}})\times\operatorname{CF}({\mathbb{R}})$. It is clear that $\nu_{r}(K,L)\in\Gamma(\text{ECT}(K)_{*}\mu,\text{ECT}(L)_{*}\mu)$ for all $r\in O(d)$. Then we have the following inequality:
$W_{1}(\text{ECT}(K)_{*}\mu$ $,\text{ECT}(L)_{*}\mu)\leq\inf_{r\in O(d)}\left(\int_{\text{Im}(\text{ECT}(K))\times\text{Im}(\text{ECT}(L))}\|f-g\|_{1}\,\mathrm{d}\nu_{r}(K,L)(f,g)\right)$
$=\inf_{r\in O(d)}\left(\int_{S^{d-1}}\|\text{ECT}(K,v)-\text{ECT}(L,r(v))\|_{1}\,\mathrm{d}\mu(v)\right)$
$\leq\inf_{r\in O(d)}d_{\text{ECT}}^{1}(K,r(L)).$
This theorem not only shows that the pushforward measure is stable under Wasserstein distance but also more stable than normal ECT outputs, as it is always less than or equal to the minimal distance between two shapes up to $O(d)$ actions.
SampEuler
The ECT pushforward measure captures all the information from shape classes of strongly generic complexes and distinguishes different shape classes. However, being a measure on a space of curves, it is hardly computable. Based on this idea, we propose the following discretization of it, called SampEuler, as a computable descriptor of shape classes.
Definition S14.
Let $K\subset{\mathbb{R}}^{d}$ be a geometric simplicial complex and let $\text{ECT}(K)_{*}\mu$ be the ECT pushforward measure of $K$. Randomly sample $(X_{1},\dots,X_{n})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\text{ECT}(K)_{*}\mu.$ The SampEuler of $K$ of order $n$, denoted $\text{SampEuler}_{n}$, is defined by
$\text{SampEuler}_{n}(K)\;=\;\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}},$
where $\delta_{x}$ is the Dirac measure at $x$.
In practice, ECT of a given geometric simplicial complex is computed by fixing a set of evenly sampled directions around the sphere along a chosen orientation. The computation of SampEuler is simply done by randomly sampling directions from the sphere for each complex and recording the corresponding ECC curve in this random order.
For a geometric simplicial complex $K$, both $\text{ECT}(K)_{*}\mu$ and $\text{SampEuler}_{n}(K)$ are measures on $\operatorname{CF}({\mathbb{R}})$. We show in Theorem S8 that $\text{SampEuler}_{n}$ converges to the ECT pushforward measure under the Wasserstein metrics as the order $n$ goes to infinity. Therefore, SampEuler captures sufficient information about the shape class $[K]$ and would be more stable than normal ECT discretizations.
We will make use of the following lemma in the proof, proven in [22] along the proof of Theorem S3.
Lemma S1.
For any simplicial complex $K$ embedded in ${\mathbb{R}}^{d}$, the support of $\text{ECT}(K)_{*}\mu$ is the set $\text{Im}(\text{ECT}(K))$.
The proof of our convergence theorem is a consequence of the following classical result.
Theorem S8 (Convergence of empirical measure [64, 65]).
Suppose $(M,d)$ is a separable metric space and $X_{1},\dots,X_{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mu\in\mathcal{P}_{r}(M)$ with $r\geq 1$. Then,
$W_{r}(\mu_{n},\mu)\rightarrow_{a.s.}0\text{ as $n\rightarrow\infty$ }.$
Theorem S9.
Given a geometric simplicial complex $K\subset{\mathbb{R}}^{d}$, denote $\mu_{n}=\text{SampEuler}_{n}(K)$ and $\mu_{K}=\text{ECT}(K)_{*}\mu$. Then, $W_{r}(\mu_{n},\mu_{K})\rightarrow_{a.s.}0\text{ as $n\rightarrow\infty$ }$.
Proof.
[…]
$\int_{\text{CF}({\mathbb{R}})}\|f-f_{0}\|_{r}\text{d}\mu_{K}(f)=\int_{\text{Im}(\text{ECT}(K))}\|f-f_{0}\|_{r}\text{d}\mu_{K}(f)\leq\int_{\text{Im}(\text{ECT}(K))}(2C))^{r}\text{d}\mu_{K}(f)=(2C)^{r},$
as $\mu_{K}$ is a probability measure. Since $\mu_{n}$ has finite support, it is trivial that $\mu_{n}\in\mathcal{P}_{r}(\operatorname{CF}({\mathbb{R}}))$
Next, we show that the space of possible Euler Characteristic Curves (ECCs) is separable. We know from Theorem S1 that this space is a Hausdorff metric space using $L_{p}$ metrics with $1\leq p<\infty$. Take curves that change values at n rational numbers; this set $F_{n}$ is in one-to-one correspondence with the set $({\mathbb{Q}}\times{\mathbb{Z}})^{n}$, which is countable. Take the countable union $F=\cup_{i=0}^{\infty}F_{n}$. Since each constructible function $f\in\text{CF}({\mathbb{R}})$ only has finitely many value changes, we can find $f^{\prime}\in F$ arbitrarily close to $f$. Since $F$ is countable, the space $\text{CF}({\mathbb{R}})$ is separable. ∎
Vectorization of SampEuler
As shown above, computing ECT along random directions using SampEuler has nice theoretical guarantees. However, for small biological datasets, this would lead to a high-dimensional feature, and the classification models would suffer from the curse of dimensionality. Interpreting the pushforward measure result also requires some visualization tools to help. To tackle the two problems, we propose the following vectorization algorithm, which can be used as a low-dimensional vectorization tool as well as a visualization tool of the SampEuler. It evaluates the empirical ECT pushforward measure, computed by SampEuler, on the sets of curves having constant value across certain intervals.
Definition S15.
Let $K$ be an embedded simplicial complex in ${\mathbb{R}}^{d}$ and $l>0$. Denote by $\text{Int}({\mathbb{R}})_{l}$ the set of closed intervals of length $l$ and $K_{(a,b),c}:=\{\varphi\in\text{supp}(\text{SampEuler}_{n}(K)):\varphi|_{(a,b)}=c\mathbf{1}_{(a,b)}\}$ for $a,b\in{\mathbb{R}}$ and $c\in{\mathbb{Z}}$. We define the vectorization of SampEuler(or vectorized SampEuler) of $K$ as follows:
$\mathcal{V}(K,l):\text{Int}({\mathbb{R}})_{l}\times{\mathbb{Z}}$ $\rightarrow[0,1],$
$([a,b],k)$ $\mapsto\text{SampEuler}_{n}(K)(K_{[a,b],k}),$
We will often abbreviate the notation to be $\mathcal{V}(K)([a,b],k)$ when $l$ is clear from the context.
Note that the above definition is still well-defined when we use the ECT pushforward measure instead of the SampEuler of $K$. For theoretical purposes, we work with the definition using ECT pushforward measures. It is clear that if two complexes are equal up to isometry, they have the same ECT pushforward measure. Hence, the evaluations of this vectorization for both complexes are equal when evaluated using the ECT pushforward measure. This shows that such vectorization is isometry invariant.
Remark 1.
We can recover the average of ECT over all directions at given filtration values using the vectorization of SampEuler by:
$\int_{S^{d-1}}\text{ECT}(K)(v,t)\,\mathrm{d}v=\int_{{\mathbb{R}}}\mathcal{V}(K,0)({t},y)\,\mathrm{d}y.$
Using this observation, we can write down the explicit formula for computing DETECT [28] from the vectorization of SampEuler:
$\text{DETECT}(K)(x)$ $=\int_{S^{d-1}}\int_{-a}^{x}\text{ECT}(K)(v,t)-\overline{\text{ECT}(K)(v,-)}\,\mathrm{d}t\,\mathrm{d}v$
$=\int_{-a}^{x}\int_{S^{d-1}}\text{ECT}(K)(v,t)\,\mathrm{d}v\,\mathrm{d}t-\int_{-a}^{x}\int_{S^{d-1}}\overline{\text{ECT}(K)(v,-)}\,\mathrm{d}v\,\mathrm{d}t$
$=\int_{-a}^{x}\int_{-\infty}^{\infty}\mathcal{V}(K,0)(t,y)\,\mathrm{d}y\,\mathrm{d}t-\left(\frac{x+a}{2a}\right)\int_{S^{d-1}}\int_{-a}^{a}\text{ECT}(K)(v,t)\,\mathrm{d}t\,\mathrm{d}v$
$=\int_{-a}^{x}\int_{-\infty}^{\infty}\mathcal{V}(K,0)(t,y)\,\mathrm{d}y\,\mathrm{d}t-\left(\frac{x+a}{2a}\right)\int_{-a}^{a}\int_{-\infty}^{\infty}\mathcal{V}(K,0)(t,y)\,\mathrm{d}y\,\mathrm{d}t.$
This shows that vectorization of SampEuler contains all the information from DETECT while still being isometry invariant.
Computation Implementations
In this section, we discuss how ECT, SampEuler, and the vectorization of SampEuler are computed in applications for geometric simplicial complexes and grayscale images. We adopt Lemma S2 from [22].
Definition S16.
Let $K$ be a geometric simplicial complex. For each vertex $x\in K$, the lower star of $x$ along direction $v$, denoted by $K^{(x,v)}$, is defined by:
$K^{(x,v)}:=\{\sigma\subset K:\max_{y\in\sigma}y\cdot v=x\cdot v\}.$
Lemma S2.
Let $K$ be a geometric simplicial complex. Denote by $\text{ver}(K)$ the set of vertices of $K$. Then we have:
$\text{ECT}(K)(v,-)=\sum_{x\in\text{ver}(K)}\chi(K^{(x,v)})\mathbf{1}_{[x\cdot v,\infty)}.$
Using Lemma S2, we compute the ECT of $K$ as follows. We first fix a set of directions along a chosen orientation. For each chosen direction $v$, we calculate the vertex of maximum filtration value of each $k$-simplex $x_{0},\dots,x_{k}]$, i.e. we find $x_{*}=\text{argmax}\{x_{i}\cdot v:i=0,\dots,k\}$. We assign this $k$-simplex to the lower star of $x_{*}$. We compute the Euler characteristic of the lower star of all vertices after assigning all simplices. Without loss of generality, we assume the filtration values of all vertices are bounded in the interval [a,a][-a,a]. We densely sample $t$-values in the interval [a,a][-a,a]. For each $t$, we compute the sum $\sum_{\begin{subarray}{c}x\in\text{ver}(K)\\ x\cdot v\leq t\end{subarray}\chi(K^{(x,v)})}$ as the value of $\text{ECT}(K)(v,t)$. We record all the outputs in order, and end up with an $m\times n$ matrix representing ECT, where $n$ is the number of directions computed and $m$ is the number of $t$-values evaluated along each direction. See Figure S2 for illustration. An illustration video can be found in https://github.com/reddevil0623/Shape-transform-descriptor-for-thymus-structures. The computation of ECT for grayscale images is done similarly, and see [45] for more details.
The computation of SampEuler is similar, but with a random set of directions generated instead of the fixed set of directions. We now explain the computation of the vectorization based on the results of SampEuler. Without loss of generality, we assume the filtration values of all vertices along all directions are bounded within the interval [a,a][-a,a] and the Euler characteristics of all possible sublevel sets are bounded within the interval [b,b][-b,b]. Let ${\mathbb{Z}}_{b}={\mathbb{Z}}\cap[-b,b]$. We uniformly partition the space $a,a]\times{\mathbb{Z}}_{b}$ into constant intervals of the same length. We consider the matrix representation of SampEuler. For each curve represented as an array of numbers $c_{1},\dots,c_{m}]$, let $c_{i},\dots,c_{j}$ be the entries with corresponding filtration value in the interval $l$. If $c_{i},\dots,c_{j}$ are all equal to a constant $c$, we add 1 to the corresponding entry in the vectorization result for interval $l\times c$. Otherwise, nothing changes. We repeat the process for all curves and divide the resulting matrix by $n$, the number of directions, to normalize.
Figure S2: A: Schematic showing how the Euler characteristic transform (ECT) is calculated for complex $K$ along direction $v$ at filtration value t. The hyperplane $x\cdot v=t$ is drawn to determine the half-space $x\cdot v<t$. The vertices of $K$ lying in this half space, together with their lower stars, are considered as plotted in red. The Euler characteristic of the red part of the shape is then computed as the output $\text{ECT}(K)(v,t)$. B: A segmented binary quadrant of mouse thymi. The image is of size 200 pixels $\times$ 200 pixels. We use [45] to create a square complex of diagonal length 1 based on the image for computing ECT. C: Euler characteristic curve along direction $(1,0)$ for the image in B. The initial increase in the curve value accounts for the emergence of disjoint white pixels on the left edge of the image. As we include more of the image towards the right, white pixels connect and form loops, aligning with the fast decrease starting around filtration value 0.25-0.25. The final increasing trend at around filtration value $0.25$ is again due to the increase in connected components on the right end of the image, due to the cutting of the quadrant.
S2B Other Methods
Persistence Homology with Signed Distance Function
We use [66] to compute the signed distance function filtration for each thymus image by taking the boundary as the boundary contour of the white pixels. We apply [37] to compute degree 0 and degree 1 Persistence Image for each image using signed distance function filtration with a Gaussian kernel $\sigma=1.0$ and uniform weight.
Enrichment ratio
To characterize changes in cell-type composition within each thymic region, a naive approach is to examine the raw proportion of each cell type in that region. However, with more than 40 cell types, Presel DP comprises around $60\%$ in both age groups. This makes changes in the remaining types be obscured when using proportions alone. To better capture meaningful differences across all cell types, we instead use a cell enrichment ratio, defined as follows.
Definition S17.
Consider quadrant $q\in Q$, and cell type $j\in J$. Let $C_{q,j}$ be the number of type $j$ cells in quadrant $q$. Denote
$r_{q,j}=\frac{C_{q,j}+\alpha}{\sum_{k\in J}C_{q,k}+\alpha J},$
where $\alpha$ is the Dirichlet smooth constant set to avoid quadrants with few cells, and $r_{q,j}$ is the smoothed ratio of type $j$ cell in quadrant $q$. Similarly, we denote
$Y_{j}=\frac{(\sum_{q^{\prime}\in Y}C_{q^{\prime},j})+\alpha}{(\sum_{k\in J}\sum_{q^{\prime}\in Y}C_{q^{\prime},j})+\alpha J}$
as the ratio of type $j$ cells in all young quadrants to be used as the baseline. We define the enrichment ratio for quadrant $q$ and cell type $j$ as
$E_{q,j}=\frac{r_{q,j}}{Y_{j}}.$
We focus on cell types: DN3, Postsel DP(CD69+), Postsel DP(CD69-), Presel DP. For each quadrant $q$, we have a cell feature vector $(E_{q,1},\dots,E_{q,n})$. We equip the cell feature vector space with the $L_{2}$ Euclidean norm.
Appendix S3 Data Analysis
S3A SampEuler Examples
In this section, we show how vectorization of SampEuler provides interpretable topological features. The vectorization creates an intensity map of ECC curves across all directions, revealing typical topological patterns for a given shape. We demonstrate how these patterns can be interpreted in terms of the underlying geometric structure.
We generate sets of simplicial complex representations of disjoint ellipses of major axis length 2 and minor axis length 1. We sample 80 points around the boundary of each ellipse and apply Gaussian noise with standard deviation $0.05$. We create $n$ such ellipses by randomly sampling the position of the centre and the angle of rotation. We fill all ellipses. See Figure S3.
We first show that the vectorized SampEuler can detect the connectedness of the input complex. In Figure S3A-B, the ellipse centres are uniformly distributed in a square of side length 50. Figure S3A contains 50 such ellipses, and Figure S3B contains 40 such ellipses. By definition, the final converging value of the vectorized SampEuler is equal to the total Euler characteristic of the input complex. In this case, the Euler characteristic is equal to the number of connected components. The converging value is 50 in Figure S3A and 40 in Figure S3B, indicating that Figure S3A has more connected components.
In Figure S3C-D, we show that vectorized SampEuler captures different sampling distributions used for sampling ellipse centres. The centres are restricted to quadrants 2, 3, and 4 for Figure S3C. Two general trends appear in the Euler characteristic curves. The first trend begins near filtration value 30-30 and rises to a peak near +10+10. This corresponds to directions pointing into quadrant 1: the growing sublevel set first encounters the ellipses at filtration values comparable to those in Figure S3A. However, the final ellipse has a lower filtration value than in Figure S3A due to the missing corner in the first quadrant, and the ellipses are more densely packed in Figure S3C, explaining the steeper trend of the curve. The second trend starts around 20-20 and peaks at about +20+20. This reflects curves along directions in the third quadrant: in the absence of ellipses in the first quadrant, the first ellipse is included later than in Figure S3A, shifting both the start and peak to higher filtration values. The ending filtration value matches because both Figure S3C and Figure S3A include their final ellipses, the corner ellipse in the third quadrant, at similar heights. For all other directional curves, the increasing section lies between these two extremes, yielding the “quadrilateral” region in the vectorized SampEuler plot.
The vectorized SampEuler in Figure S3D displays a funnel-shaped region bounded approximately by lines connecting (40,0)(-40,0) to $(40,50)$ and (10,0)(-10,0) to $(10,50)$. This pattern reflects the geometry of the elliptical sampling region. For directions near the minor axis, ellipses first appear at higher filtration values but are captured within a short interval, producing steep ECC curves. For directions near the major axis, ellipses appear earlier but span a longer filtration range, producing shallower curves. The funnel shape emerges as ECCs transition continuously between these extremes across all sampling directions.
Figure S3: A: i) 50 geometric simplicial complex representations of ellipses with centres uniformly sampled from the centred square of side length 50. ii) The corresponding vectorized SampEuler result. B: i) 40 geometric simplicial complex representations of ellipses with centres uniformly sampled from the centred square of side length 50. ii) The corresponding vectorized SampEuler result. C: i) 50 geometric simplicial complex representations of ellipses with centres uniformly sampled from the restriction of quadrant 2, 3, 4 of the centred square of side length 50. ii) The corresponding vectorized SampEuler result. D: 50 geometric simplicial complex representations of ellipses with centres uniformly sampled from a centred ellipse with major axis length 100 and minor axis length 20. ii) The corresponding vectorized SampEuler result.
S3B Comparison to ECT and DETECT
In this section, we compare SampEuler and vectorized SampEuler to established topological shape descriptors. This example demonstrates that SampEuler and vectorized SampEuler capture sufficient information about the shape class of the input complex.
We use geometric simplicial complexes explained in Section S1A to illustrate that SampEuler and its vectorization are more robust to perturbations than the ECT, and are more informative than DETECT. We run two experiments for each method. In the first experiment, we generate two families of 20 complexes, with each complex randomly rotated about the origin. In the second experiment, we generate complexes without rotating them. We use mdscale [67] to visualize the results. As shown in Figure S4, ECT only distinguishes the two families of shapes correctly when all complexes are aligned. However, when we randomly rotate the complexes, the ECT fails to distinguish the two families. For DETECT, information loss during averaging means that DETECT fails to distinguish the two families in both experiments. However, with SampEuler and vectorized SampEuler, clear separation of the two classes can be seen in both experiments, showing that they are robust to isometries and contain sufficient information about the shape of input complexes.
Figure S4: MDS visualizations of pairwise similarities between analysis results of two classes of generated geometric simplicial complexes. The left plots are when all complexes are aligned, and the right plots are when all complexes are randomly rotated. A: results of ECT. B: results of DETECT. C: results of SampEuler. D: results of vectorized SampEuler.
S3C MPEG-7 Dataset for Shape Classification
To test against existing methods and understand the impact of discretization on analysis results, we apply our methods to the MPEG-7 dataset with simple normalization steps, as described in Section S1B. We first compare with methods mentioned in [33, 44]. We use the 70/30 train-test-split ratio and repeat the experiment 100 times to report the average accuracy. To be consistent with other topological pipelines, we use the Support Vector Machine (SVM) classification model [68]. We compute the pairwise Wasserstein distance for SampEuler and $L_{2}$-distances for vectorized SampEulers and construct the RBF-kernel as input to the SVMs. Classification accuracy is reported using the best-performing parameters from the ablation study (Section S3D).
Table S1: The average accuracy $(\%)$ of different methods on the 10-class subset of the MPEG7 dataset.
Method Accuracy
Persistence Scale Space Kernel [29] 73.33
Persistence Weighted Gaussian Kernel [30] 74.83
Sliced Wasserstein Kernel [31] 76.83
Tangent Vector Representation with Gaussian Kernel [32] 66.17
Persistence Fisher Kernel [33] 80.00
Tree Sliced Wasserstein Kernel [44] (80.00, 85.00)*
QUPID [34] 68.80
ECT [17] 86.23
DETECT [28] 51.02
SampEuler 91.67
vectorized SampEuler 94.39
* Exact value not reported; range estimated from the bar chart in the original text.
To compare with other dedicated 2D pattern recognition methods and topological methods with deep learning, we also perform the classification task on the whole MPEG7 dataset with a 90/10 train-test-split ratio and repeat the experiments 10 times to report the average accuracy.
Table S2: The average accuracy (%) of different methods on the MPEG7 dataset.
Method Accuracy
Skeleton Paths [35] 86.7
Bag of Contours [36] 97.2
Persistence Images + NN [37] 92.3
Learned Persistence Codebooks [38] + NN 92.7
ECT [17] 71.3
DETECT [28] 36.4
SampEuler 91.4
vectorized SampEuler 86.8
X indicates our proposed methods.
As shown in LABEL:table:_accuracy_1, SampEuler and vectorized SampEuler provide better classification accuracies than all other methods, with vectorized SampEuler attaining the highest accuracy, likely because its vectorization step reduces feature dimensionality while maintaining all shape information. In the larger dataset of LABEL:table:_accuracy_2, Bag-of-Contours and neural-network–based methods attain higher accuracies than our approaches. The Bag-of-Contours method in [36] is tailored to 2D shape recognition and relies on precise boundary-contour segmentation, which can be computationally intensive and dataset-dependent. Likewise, neural networks often achieve state-of-the-art accuracy on sufficiently large training sets, though their learned representations are uninterpretable and demand substantial computational resources. Although SampEuler and vectorized SampEuler fall short of these top accuracies, they are segmentation-free, applicable to general shape analysis, and come with theoretical guarantees that support interpretability. In this larger dataset, the original SampEuler achieves better performance than its vectorized variant, presumably because its higher-resolution features can leverage the increased sample size to capture finer geometric details.
Across both datasets, SampEuler and vectorized SampEuler consistently outperform ECT- and DETECT-based baselines. DETECT suffers substantial information loss from the averaging step, leading to the worst performance overall. The high sensitivity of ECT to perturbations causes instability in classification tasks.
S3D Ablation Study
We use the MPEG7 dataset to perform the ablation study of the SampEuler method. The results are shown in Figure S5 and Figure S6. These results suggest that if we sample more than 100 directions and 1000 points along each direction, the additional information yields diminishing returns in classification accuracy. Moreover, leveraging the optimised implementation in Eucalc [45] enables efficient computation of these descriptors, making the above parameter choices practical in routine use.
Refer to caption $xpoint=100$ Refer to caption $xpoint=1000$ Refer to caption $xpoint=3000$ Refer to caption $xpoint=5000$ Refer to caption $xpoint=7000$ Refer to caption $xpoint=9000$ Refer to caption $xpoint=11000$ Refer to caption $xpoint=13000$ Refer to caption $xpoint=15000$
Figure S5: Ablation study of SampEuler by fixing the number of directions sampled and varying the number of filtration values along each direction.
Refer to caption $k=10$ Refer to caption $k=100$ Refer to caption $k=200$ Refer to caption $k=300$ Refer to caption $k=400$ Refer to caption $k=500$ Refer to caption $k=600$ Refer to caption $k=700$
Figure S6: Ablation study of the SampEuler by fixing the number of filtration values along each direction and varying the number of directions sampled.
S3E Quantifying Thymus Morphological Difference
We report mean classification accuracy across 50 trials and the wall-clock runtime of the ECT-based pipelines. For ECT-based methods, we compute 360 directions with 300 points along each direction. SampEulers are computed using the same parameters but with random set of directions. As a baseline, we benchmark against persistence-diagram (PD) features [37] computed from signed-distance-function (SDF) filtrations [66]. Concretely, for each binary image we compute its SDF, construct the sublevel-set filtration, and extract PDs in homology dimensions 0 and 1, thereby capturing births and deaths of connected components and loops while encoding geometry via distance to the boundary. We determine global birth and persistence ranges to set the bounding boxes, discard points with infinite death, and discretize Persistence Images (PIs) with a pixel size of $\tfrac{\text{global persistence range}}{500}$. PIs use a Gaussian kernel with $\sigma=1.0$ and uniform weights; the images are then flattened into feature vectors. For both the ECT-based features and PIs, we train a random-forest classifier: we select hyperparameters via a fixed grid, refit with the best setting, and evaluate across trials.
We also compare against state-of-the-art deep learning models for image classification [46, 47, 48, 49]. For AlexNet, ViT, MobileNetV2, and EfficientNet-B0, we initialize from ImageNet-pretrained weights and replace the final layer with a two-way linear head. Inputs are resized to $224\times 224$ RGB and normalized with ImageNet statistics. In each trial, we train using stochastic gradient descent (learning rate $10^{-3}$, momentum $0.9$), batch size $16$, and cross-entropy loss for $5$ epochs.
Table S3: The average accuracy $(\%)$ and run time of different methods on thymus K8 quadrants and K14 quadrants.
Method K8 Acc. K14 Acc. K8 Time K14 Time
[…]
We use established methods to confirm our interpretations of the thymus structures. We compute the zeroth and first Betti numbers of images for the number of connected components and loops. The results are plotted as violin plots in Figure 7(a). For K8, the violin plots for the zeroth Betti number show that young quadrants have a slightly higher number of connected components. Similarly, Betti-1 plots show that old quadrants have a higher rank for $H_{1}$, meaning that old quadrants contain more loops. For K14, the violin plots for Betti-0 and Betti-1 both show that the old quadrants have a much larger range of variation compared with the young quadrants. These observations are consistent with our interpretations of the SHAP scores of vectorized SampEuler features.
To assess the density of holes in the cortex component, we apply the signed distance function filtration to each image so that the centres of the holes have negative distances. Now, the persistence of zeroth-degree features in persistent homology represents how dense the holes are. The further away the holes are, the longer these zeroth-degree features will persist. As shown in Figure 7(b), the young quadrants have more persistent characteristics than the old quadrants. These observations are consistent with our interpretation from the gradient of vectorized SampEuler computation.
(a) Violin plots of Betti numbers 0 & 1.
(b) Difference of persistence images: old – young.
Figure S7: (a) Violin plots of Betti number 0 and Betti number 1 from top 5% to 95% of samples. (b) Difference of persistence images: $\mathrm{P}_{\mathrm{old}}-\mathrm{P}_{\mathrm{young}}$.
S3F Depth Analysis
We apply depth analysis of the cortical region to preserve the spatial heterogeneity. Since the thymus and medulla boundaries are not homotopic, we choose the following definition as the notion of depth.
Definition S18.
Given a pixel $p$ in the cortical region of the thymus, we define:
$d_{p}:=\min\{d(p,x),\text{ for }x\text{ in medulla}\}.$
where $d$ is the usual Euclidean distance. For a quadrant $Q$ in thymus $T$, we define:
$d_{Q}:=\frac{1}{|Q|}\sum_{p\in Q}d_{p}.$
We define the normalized depth of $Q$ as follows:
$D(Q):=1-\frac{d_{Q}}{\max_{Q^{\prime}\in T}d_{Q^{\prime}}}.$
Remark 2.
Since the medulla region consists of more than one connected component in all thymic samples, there is no homotopy between the thymic capsule and the cortex-medulla junction (CMJ). The above notion of depth is used when the CMJ is of interest to the study.
One could also define the distance $d_{p}$ as the minimal distance to any thymic boundary point and obtain another notion of normalized depth with 0 at the capsule and 1 at the centre of the thymus. This notion of depth is more suitable when the whole thymus is of interest. We use it later for the DN cell analysis.
Based on the normalized depth of a quadrant, we construct the following depth kernel for each quadrant:
$w_{Q}(t)=\exp\left(-\frac{1}{2}\left(\frac{D(Q)-t}{h}\right)^{2}\right).$
We use $h=0.1$ and normalize the weight by each age group $A$ and get the normalized kernel $\tilde{w}_{Q}(t)=\frac{w_{Q}(t)}{\sum_{Q^{\prime}\in A}w_{Q^{\prime}}(t)}$. To check that the depth notion is consistent for both age groups, we plot the distribution of quadrants at various depth intervals in Figure S8. The result suggests that the shrinkage in size of the thymus due to ageing does not affect our notion of depth, and so it is suitable to compare quadrants at the same depth from different age groups to preserve the spatial heterogeneity of the cortical region.
Figure S8: Depth distributions of quadrants of each age group. The x-axis is the normalized depth $D(Q)$ of each quadrant $Q$, and the y-axis is the count of quadrants having depth in a given depth interval.
For depth $t$, the normalized kernels define probability measures on quadrants for each age group,
$P_{Y}^{t}$ $=\sum_{Q\in Y}\tilde{w}_{Q}^{Y}(t)\,\delta_{Q},$ (S1)
$P_{O}^{t}$ $=\sum_{Q\in O}\tilde{w}_{Q}^{O}(t)\,\delta_{Q},$ (S2)
where $\delta_{Q}$ is the Dirac measure of the corresponding feature of $Q$. We can then compute the energy distance [50] $E(P_{Y}^{t},P_{O}^{t})$ between the two measures using the pairwise distance matrix $M$ of quadrant features by
$E(P_{Y}^{t},P_{O}^{t})$ $=2\left(\sum_{i\in Y}\sum_{j\in O}\tilde{w}_{i}^{Y}(t)\tilde{w}_{j}^{O}(t)M_{ij}\right)$ (S3)
$\quad-\sum_{i,j\in O}\tilde{w}_{i}^{O}(t)\tilde{w}_{j}^{O}(t)M_{ij}-\sum_{i,j\in Y}\tilde{w}_{i}^{Y}(t)\tilde{w}_{j}^{Y}(t)M_{ij}.$
S3G Cell Distribution vs Morphology
For the same set of mouse thymi, we obtained cell‐type and spatial locations for all cells within each thymus. Our goal is to characterize and compare age‐related changes within the thymus and the epithelial architecture of the thymus. To this end, we perform k-medoids clustering [51] separately on cell-location features and on shape descriptors of thymic quadrants, and then compare the resulting cluster partitions.
To allow faster computation of the pairwise Wasserstein distance matrix while keeping sufficient information, we sample $100$ directions with $3000$ points along each direction, and we use the sliced Wasserstein distance algorithm [53] with $50$ slices.
Since we focus on the cortical compartment, and for simplicity, we restrict the compositional analysis of haematopoietic cells to DN3, Presel DP, Postsel DP (CD69+) and Postsel DP (CD69-) thymocytes. These cell types not only constitute the vast majority of cortical thymocytes but are also centered around the developmentally-relevant process of positive selection.
We discuss how to choose the number of clusters for both clusterings. We first applied silhouette analysis [51] to find the most suitable number of clusters from the set $\{2,3,\dots,10\}$ for both k-medoids clusterings. The analysis suggests $k=2$ for both the clustering based on enrichment ratios and the clustering based on SampEulers. The results are shown in Figure S9 and Figure S10. Two clusters are not able to provide meaningful biological interpretations about the impact of the ageing process. Therefore, we manually force three clusters for both clusterings as the second-highest silhouette score choice. To further constrain the analysis to look at cells in direct contact with thymic epithelial cells (TECs), we filter to only count cells of selected cell types that are at a distance less than or equal to $5\text{ }\mu\text{m}$ from a TEC. The results are shown in Figure S11 and Figure S12.
Figure S9: K-medoids clustering results with 2 clusters for all thymic cortex quadrants. This plot contains result for the first four thymi, the remaining results are included in Figure S10. The first column on the left is the results using SampEulers, the second column is the results using cell enrichment ratios.
Figure S10: K-medoids clustering results with 2 clusters for all thymic cortex quadrants. This plot contains result for the remaining thymi continuing from Figure S9. The first column on the left is the results using SampEulers, the second column is the results using cell enrichment ratios.
Figure S11: K-medoids clustering results for all thymic cortex quadrants using 3 clusters. This plot contains results for the first four thymi, the remaining results are included in Figure S12. The first column on the left is the results using SampEulers, the second column is the results using selected cell enrichment ratios.
Figure S12: K-medoids clustering results for all thymic cortex quadrants using 3 clusters. This plot contains results for the remaining thymi continuing from Figure S11. The first column on the left is the results using SampEulers, the second column is the results using cell enrichment ratios.
[…]
Table S4: PhenoCycler antibody panel with host species, clone, vendor, catalogue number, fluorophore, dilution, and exposure time. DAPI was imaged at every round. For the other markers, the order in which they appear in the table represents the order in which they were imaged in groups of three (see Materials and Methods). Ar Hms = Armenian hamster, Ms = mouse, Rb = rabbit, Rt = rat, Sy Hms = Syrian hamster
Marker Host species Clone Vendor Catalogue number Fluorophore Dilution Exposure time [ms]
DAPI - - - - - 1:1000 150
AIRE Rt 5H12 Invitrogen 14-5934-82 Cy5 1:200 200
B220 Rt Ra3-6B2 Akoya 4150006 AF488 1:200 200
CD8A Rt 53-6.7 Akoya 4250017 ATTO550 1:200 150
FOXP3 Rt FJK-16s Invitrogen 14-5773-82 Cy5 1:200 200
TCRGD Ar Hms GL3 BioLegend 118128 AF488 1:100 300
NP63 Rb polyclonal Proteintech 12143-1-AP ATTO550 1:200 300
PD1 Rt 29F.1A12 Invitrogen P378-1MG Cy5 1:200 300
CD86 Rt GL-1 Invitrogen MA1-10299 AF488 1:200 300
CD69 Ar Hms H1.2F3 Invitrogen 14-0691-82 ATTO550 1:200 300
CD11C Ar Hms N418 Akoya 4350013 Cy5 1:200 300
CD45 Rt 30-F11 Akoya 4150002 AF488 1:200 150
ICOS Ar Hms C398.4A Invitrogen 14-9949-82 ATTO550 1:200 200
XCR1 Ms ZET BioLegend 148202 Cy5 1:100 300
PDCA1 Rt eBio927 Invitrogen 16-3172-81 AF488 1:200 300
SIRPA Rt P84 Invitrogen 16-1721-81 ATTO550 1:100 300
LY6G Rt 1A8 Akoya 4350015 Cy5 1:200 100
CD71 Rt R17217 Invitrogen 17-0711-82 AF488 1:200 300
KI67 Ms B56 Akoya 4250019 ATTO550 1:200 150
SCA1 Rt D7 Invitrogen 14-5981-85 Cy5 1:200 200
CD25 Rt PC61.5 Invitrogen 14-0251-86 AF488 1:200 300
HELIOS Ar Hms 22F6 BioLegend 137202 ATTO550 1:200 200
K10 Rb polyclonal Invitrogen PA5-104456 Cy5 1:400 80
CD62L Rt 95218 R&D MAB5761 AF488 1:200 200
CD5 Rt 53-7.3 Akoya 4250007 ATTO550 1:200 250
ERTR7 Rt ER-TR7 Bio-Rad MCA2402 Cy5 1:400 80
CD24 Rt M1/69 Akoya 4150014 AF488 1:200 200
CD31 Rt MEC13.3 Akoya 4250001 ATTO550 1:200 200
CD49F Rt GoH3 Akoya 4350007 Cy5 1:200 200
CD11B Rt M1/70 Akoya 4150015 AF488 1:200 200
CD4 Rt RM4-5 Akoya 4250016 ATTO550 1:200 150
LY51 Rt 6C3 BioLegend 108302 Cy5 1:100 300
PIGR Rb polyclonal Invitrogen PA5-35340 AF488 1:200 250
MHCII Rt M5/114.15.2 Akoya 4250003 ATTO550 1:400 150
TCRB Ar Hms H57-597 Akoya 4350006 Cy5 1:200 150
PDPN Sy Hms 8.1.1 Invitrogen MA5-18054 AF488 1:200 300
CD206 Rt C068C2 BioLegend 141701 ATTO550 1:400 250
K8 Rt TROMA-1 EMD Millipore MABT329 Cy5 1:200 80
K14 Rb polyclonal Invitrogen PA5-16722 AF488 1:200 80
CD44 Rt IM7 Akoya 4250002 ATTO550 1:200 200
Table S5: List of PhenoCycler cell phenotypes with their defining marker thresholds, distance thresholds, and regional filters.
No. Phenotype name Corresponding cell type Defining markers and their normalized expression thresholds Distance thresholds [pixels] Region filter
1 ICOS+ T-cells expressing ICOS ICOS > 0.25 7 -
2 T-reg Regulatory T-cells FOXP3 > 0.15 6 -
3 CD11b+ Monocytes CD11b > 0.15 8 -
4 T-gd (PD1+) gd T-cells expressing PD1 (TCRgd > 0.15) & (PD1 > 0.25) 6 -
5 T-gd (PD1-) gd T-cells not expressing PD1 (TCRgd > 0.15) & (PD1 $\leq$ 0.25) 6 -
6 DN3 DN T-cells (CD44 $\leq$ 0.2) & (CD25 > 0.2) & (CD5 $\leq$ 0.2) & (TCRb $\leq$ 0.2) 6 Cortex
7 ISP8 Immature SP8 T-cells (CD25 $\leq$ 0.2) & (CD8a > 0.3) & (CD4 $\leq$ 0.2) & (CD5 $\leq$ 0.2) & (TCRb $\leq$ 0.2) - Cortex
8 SP8 (CD69+) SP8 T-cells expressing CD69 (CD8a > 0.3) & (CD4 $\leq$ 0.2) & ((TCRb > 0.15) $\mid$ (CD5 > 0.15)) & (CD69 > 0.25) - Medulla
9 SP8 (CD69-) SP8 T-cells not expressing CD69 (CD8a > 0.3) & (CD4 $\leq$ 0.2) & ((TCRb > 0.15) $\mid$ (CD5 > 0.15)) & (CD69 $\leq$ 0.25) - Medulla
10 MP Macrophages CD206 > 0.075 8 -
11 BC B cells B220 > 0.1 4 for BC clusters, 6 for interspersed BC -
12 SP4 (Helios+) SP4 T-cells expressing Helios (CD8a $\leq$ 0.2) & (CD4 > 0.2) & ((TCRb > 0.15) $\mid$ (CD5 > 0.15)) & (Helios > 0.15) - Medulla
13 pDC (PDCA1+) Plasmacytoid dendritic cells expressing PDCA1 (CD11c > 0.25) & (PDCA1 > 0.25) 11 -
14 cDC1 (XCR1+) Conventional dendritic cells expressing XCR1 (CD11c > 0.25) & (XCR1 > 0.35) 11 -
15 cDC2 (SIRPa+) Conventional dendritic cells expressing SIRPa (CD11c > 0.25) & (PDCA1 > 0.25) 11 -
16 DC (CD11c+) Remaining dendritic cells expressing CD11c CD11c > 0.25 11 -
17 SP4 (CD69+) SP4 T-cells expressing CD69 (CD8a $\leq$ 0.2) & (CD4 > 0.2) & ((TCRb > 0.2) $\mid$ (CD5 > 0.15)) & (CD69 > 0.25) - Medulla
18 SP4 (CD69-) SP4 T-cells not expressing CD69 (CD8a $\leq$ 0.2) & (CD4 > 0.2) & ((TCRb > 0.2) $\mid$ (CD5 > 0.15)) & (CD69 $\leq$ 0.25) - Medulla
19 Postselection DP (CD69+) DP T-cells that have undergone selection and express CD69 (CD4 > 0.2) & (CD8a > 0.2) & ((CD5 > 0.3) $\mid$ (TCRb > 0.2)) & (CD69 > 0.25) - Cortex
20 Postselection DP (CD69-) DP T-cells that have undergone selection and do not express CD69 (CD4 > 0.2) & (CD8a > 0.2) & ((CD5 > 0.3) $\mid$ (TCRb > 0.3)) & (CD69 $\leq$ 0.25) - Cortex
21 Preselection DP DP T-cells that have not undergone selection (CD4 > 0.05) & (CD8a > 0.05) & ((CD5 $\leq$ 0.3) $\mid$ (TCRb $\leq$ 0.3)) & (CD69 $\leq$ 0.3) - Cortex
22 mTEC (Aire+) mTEC expressing Aire Aire > 0.09 9 -
23 mTEC (Ly6g+) mTEC expressing Ly6g Ly6g > 0.09 14 -
24 mTEC (PIGR+) mTEC expressing PIGR PIGR > 0.18 11 Medulla
25 mTEC (K10+) mTEC expressing K10 K10 > 0.1 - -
26 mTEC (CD49f+) mTEC expressing CD49f (K14 > 0.1 $\mid$ K8 > 0.1) & (CD49f > 0.3) & (PDPN < 0.15) - Medulla
27 mTEC hi (K14+, CD86+) mTEC expressing K14, CD86, and MHCII (K14 > 0.1) & (MHCII > 0.35) & (CD86 > 0.2) - -
28 mTEC hi (K8+, CD86+) mTEC expressing K8, CD86, and MHCII (K8 > 0.1) & (MHCII > 0.35) & (CD86 > 0.2) - Medulla
29 TEC (PDPN+) TEC expressing PDPN (K14 > 0.1 $\mid$ K8 > 0.1 $\mid$ Np63 > 0.1) & (PDPN > 0.15) & (ERTR7 < 0.15) 5 -
30 mTEC hi (K14+, CD86-) mTEC expressing K14, MHCII, but not CD86 (K14 > 0.1) & (MHCII > 0.35) & (CD86 $\leq$ 0.2) - -
31 mTEC hi (K8+, CD86-) mTEC expressing K8, MHCII, but not CD86 (K8 > 0.1) & (MHCII > 0.35) & (CD86 $\leq$ 0.2) - Medulla
32 mTEC lo (K14+) mTEC expressing K14 but not MHCII (K14 > 0.1) & (MHCII $\leq$ 0.35) - -
33 mTEC lo (K8+) mTEC expressing K8 but not MHCII (K8 > 0.1) & (MHCII $\leq$ 0.35) - Medulla
34 cTEC hi (Ly51+) cTEC expressing Ly51 and MHCII (K8 > 0.1) & (MHCII > 0.35) & (Ly51 > 0.2) - Cortex
35 cTEC lo (Ly51+) cTEC expressing Ly51 but not MHCII (K8 > 0.1) & (MHCII $\leq$ 0.35) & (Ly51 > 0.2) - Cortex
36 cTEC hi (Ly51-) cTEC expressing MHCII but not Ly51 (K8 > 0.1) & (MHCII > 0.35) & (Ly51 $\leq$ 0.2) - Cortex
37 cTEC lo (Ly51-) cTEC expressing neither Ly51 nor MHCII (K8 > 0.1) & (MHCII $\leq$ 0.35) & (Ly51 $\leq$ 0.2) - Cortex
38 Endo Endothelial cells CD31 > 0.2 - -
39 Fb (PDPN+, ERTR7-) Fibroblasts expressing PDPN but not ERTR7 (PDPN $\geq$ 0.15) & (ERTR7 < 0.15) & (Np63 < 0.1) - -
40 Fb (PDPN-, ERTR7+) Fibroblasts expressing ERTR7 but not PDPN (PDPN < 0.15) & (ERTR7 $\geq$ 0.15) & (Np63 < 0.1) - -
41 Fb (PDPN+, ERTR7+) Fibroblasts expressing both PDPN and ERTR7 (PDPN $\geq$ 0.15) & (ERTR7 $\geq$ 0.15) & (Np63 < 0.1) - -